SPECTROSCOPY OF 1,2-DIPHENYLETHANE-(H_{2}O)_{n} (n=1-3) CLUSTERS

\underline{JOSEPH R. GORD}, EVAN G. BUCHANAN, PATRICK S. WALSH, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette. IN 47907.

1,2-diphenylethane (DPE) is a prototypical flexible bichromophore which forms two conformers in the gas phase having close lying S_{1} and S_{2} states. Resonant ion-dip infrared (RIDIR) spectroscopy in the alkyl CH stretch region, 2800-3000 cm^{−1}, was used to assign the two observed isomers to gauche and anti structures having C_{2} and C_{2h} symmetry similar to those in butane.^{a} In the present work, the ultraviolet and infrared spectroscopy of DPE-(H_{2}O)_{n} (n=1-3) clusters has been studied to understand how the ’solvent’ water molecules bind to this molecule with two phenyl rings, and how the solvent binding perturbs the excited state behavior. RIDIR spectra in the OH stretch region (3350-3750 cm^{−1}) show that DPE-(H_{2}O)_{1} adopts a single conformation in which the water molecule bridges the two phenyl rings of the gauche conformer, forming π H-bonds with both rings. DPE-(H_{2}O)_{2} forms two structures, in which a water dimer is bound to either the gauche or anti DPE conformer. In the anti DPE-(H_{2}O)_{2} structure, the water dimer interacts primarily with one phenyl ring, disrupting the symmetry of DPE enough that the S_{1} and S_{2} origins are localized, splitting the origins by approximately 50 cm^{−1}. The development of these cluster structures with additional H_{2}O molecules, and the OH stretch spectra of the clusters in the excited state(s) will also be discussed.

aE. G. Buchanan, J. C. Dean, T. S. Zwier and E. L. Sibert, III J. Chem. Phys 2013, 138.